[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 25, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 25 13:38:01 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
There was a problem with the LX7 AWIPS workstation that was resolved with  
help from NCF.

The problem with the surface analysis over the weekend was due to an issue  
with the TAFB workstation at the backup desk. Rebooting the workstation  
resolved the problem.

HSU:
Tim Marchok is following up on the HWRF tracker problem for the pre-Seymour  
disturbance. It is similar to the tracker problem for the pre-Matthew  
disturbance where the position jumps around a lot due to weak gradients and  
vortex spin up, resulting in unrealistic transnational speeds that trigger  
the quality control flags.

TSB is still looking into the missing genesis areas in the GUI for EP  
disturbance 37.
Other items:
Work on the media room will continue this week.
NHC RFCs:
"RFC 192 to be implemented today to turn off TAFB experimental rainfall  
graphics.

RFC 205 to be implemented today to send GOES WV mosaic to WCOSS so the TCGI  
program will work for potential TCs near the coast of Africa. "

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161025/8c9a1915/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list